The car brake pedal is shown. Calculate how long the brake spring will extend, when the compressive force is F = 50 N. It is known that OA = 2 cm and OB = 24 cm, and to extend the spring by 1 cm the force 400 must be applied


1
Expert's answer
2021-03-04T11:55:05-0500
"k=\frac{F_1}{x_1}"

"x_2=\frac{F_2}{k}\\x_2=\frac{F_2}{F_1}x_1\\ x_2=\frac{50}{400}1=0.125\ cm"